OpenStudy (anonymous):

A bag contains 5 blue merbles, 3 red marbles , and 2 yellow marbles. you select a marble at random . What is P(yellow)? A) 2/8 B) 4/5 C)1/5

OpenStudy (anonymous):

We have a total of 10 marbles and 2 "correct" yellow ones. To find the probability we do:\[P=\frac{correct}{total}\]\[P(yellow)=\frac{yellow}{total}\]\[P(yellow)=\frac{2}{10}=\frac{1}{5}\]

OpenStudy (anonymous):

1/5

OpenStudy (anonymous):

thank yall